# Adversarial Bug Reports as a Security Risk in Language Model-Based Automated Program Repair

### v1

@@ 7.3. Recommendations
    Avoid Overreliance on Ensembles. While pre- and post-APR ensembles blocked 63–68% of attacks, they also raise false-positive risks, potentially harming developer trust.
@@ 7.4. Ethical Implications
Our work demonstrates that LLM-based APR systems can be manipulated into introducing malicious changes through adversarial bug reports. While this has clear security implications, our intent is to raise awareness of these risks and to enable the development of appropriate defenses. The prototype developed for this study and as well as all artifacts will be open-source, alongside a curated replication package to support further research and transparency.
We recognize the potential for abuse inherent in automated vulnerability injection. Therefore, consistent with responsible disclosure practices and guidance from prior work on LLM misuse (Greshake et al., 2023), we will notify affected stakeholders—particularly GitHub. No interaction with production systems or production repositories was performed during our experiments. All tests were conducted on forks under controlled conditions.
@@ 8. Threats to Validity
Internal Validity. Our manual annotation process involves expert judgment, which may introduce subjectivity. To mitigate this, we used a dual-annotator protocol with independent assessments and discussion-based resolution. Still, subtle vulnerabilities or security implications may have been overlooked.
External Validity. Our experiments focus on a subset of open-source projects and a single APR system (SWE-agent). While the selected repositories are diverse and representative of real-world codebases, our findings may not generalize to proprietary software or APR tools with substantially different architectures or training regimes. However, our adversarial methodology exploits structural properties shared by many modern APR systems: reliance on unvetted natural language input, automated patch generation via LLMs, and validation based primarily on test-suite outcomes. These properties are common to other state-of-the-art systems (Yang et al., 2024a; Bouzenia et al., 2024; Zhang et al., 2024; Wang et al., 2025). Unless such tools are explicitly hardened with security-aware defenses—a feature currently absent in most publicly available APR pipelines - we expect similar vulnerabilities to manifest.
Construct Validity. We define attack success in terms of patch generation, detection bypass, and the presence of malicious behavior. These metrics are practical, but may not fully capture downstream impact (e.g., exploitation feasibility or user-facing harm). Similarly, false positives in filter models or review tools could influence results.
@@ 9. Related Work
9.1. Automated Program Repair
@@
Benchmark suites like SWE-bench and SWE-bench Multimodal assess the functional correctness of LLM-generated patches (Jimenez et al., 2024; Yang et al., 2024b), but overlook whether those patches preserve security properties or introduce vulnerabilities.
While security concerns around LLMs have been raised in other domains (e.g., prompt injection, unsafe code from Copilot (Lyu et al., 2024)), such risks have not yet been systematically explored in the context of LLM-based APR.
Our work fills this gap by investigating the security risks of SE-agent-based APR systems, demonstrating that adversarial bug reports can reliably induce unsafe behavior, even in state-of-the-art LLM agents.
@@ 9.2.1. Attacks Against LLM Applications
The OWASP Top 10 for LLM Applications (Wilson and Dawson, 2023) highlights common attacks and issues occurring in deployed LLM applications. We ground our research on this established industry standard to make our results applicable and relevant for real-world APR systems.
@@
Foreshadowing our threat-model (Section 2), we decided to not employ Training Data Poisoning (LLM03) or utilize Supply Chain Vulnerabilities (LLM05), do not exploit Insecure Plugin Design (LLM07), and neither try to achieve Model Denial of Service (LLM04) nor Model Theft (LLM10).
@@ 9.2.2. Defenses for LLM-based Applications
Based on our analysis of potential attacks (Section 9.2.1), we group defenses info measures that try to prevent Prompt Injection attacks, and measures that try to avoid or minimize the impact of a successful attack.
@@ 10. Conclusion
Our empirical study shows that these adversarial inputs can systematically subvert even state-of-the-art APR pipelines: in our demonstration, over 90% of the crafted bug reports resulted in patches aligned with the attacker’s objective, including CVE reversions, vulnerability injection, and CI/CD abuse. Pre- and post-APR defenses, while helpful, often failed silently (best filter only blocked 47% issues) or demanded high levels of manual effort and scrutiny (Co Pilot supported review 58%). The fundamental asymmetry—low-cost attack versus high-cost defense—further compounds the challenge.

### v2

@@ 7.3. Recommendations
    Avoid Overreliance on Ensembles. While pre- and post-APR ensembles blocked 63-68% of attacks, they also raise false-positive risks, potentially harming developer trust.
@@
7.4. Guidance for Future Research
By highlighting a novel class of threats in the software supply chain, our work urges the community to reconsider long-held trust assumptions in automated repair pipelines.
We advocate for a security-aware design of APR systems that balances repair effectiveness with resilience to adversarial misuse. While current systems prioritize passing test suites, future work should explore APR alignment strategies that enforce project-specific security policies during synthesis preventing behavior such as reverting known CVE fixes or disabling authentication logic.
Additionally, the community requires standardized adversarial benchmarks to complement functional suites like SWE-bench; these new benchmarks would formalize red-teaming evaluations to measure an APR agent’s resilience against prompt injection and context manipulation.
Finally, given that pre-filtering remains a “cat-and-mouse” challenge, researchers should investigate improving human-in-the-loop (HITL) workflows. This includes developing forensic tools that visualize patch provenance and highlight security-critical changes, thereby reducing reviewer fatigue and enabling more effective oversight of automated contributions.
@@ 7.5. Ethical Implications
Our work demonstrates that LLM-based APR systems can be manipulated into introducing malicious changes through adversarial bug reports. While this has clear security implications, our intent is to raise awareness of these risks and to enable the development of appropriate defenses. The prototype developed for this study and as well as all artifacts are be open-source, alongside a curated replication package to support further research and transparency.
We acknowledge the abuse potential of automated vulnerability injection and, following responsible-disclosure practices and prior guidance on LLM misuse (Greshake et al., 2023), notified affected stakeholders, particularly GitHub. All experiments were conducted on forked repositories under controlled conditions, with no interaction with production systems.
@@ 8. Threats to Validity
Internal Validity. Our manual annotation process involves expert judgment, which may introduce subjectivity. To mitigate this, we used a dual-annotator protocol with independent assessments and discussion-based resolution. Still, subtle vulnerabilities or security implications may have been overlooked. LLMs are inherently non-deterministic. Typically this threat is mitigated by fixing the sampling temperature to 0. We were not able to perform this, as deviating from the default settings of the used APR system and utilized defenses negatively impacts Construct Validity.
External Validity. Our experiments focus on a subset of open-source projects and a single APR system (SWE-agent). While the selected repositories are diverse and representative of real-world codebases, our findings may not generalize to proprietary software or APR tools with substantially different architectures or training regimes. However, our adversarial methodology exploits structural properties shared by many modern APR systems: reliance on unvetted natural language input, automated patch generation via LLMs, and validation based primarily on test-suite outcomes. These properties are common to other state-of-the-art systems (Yang et al., 2024a; Bouzenia et al., 2024; Zhang et al., 2024; Wang et al., 2025). Unless such tools are explicitly hardened with security-aware defenses–a feature currently absent in most publicly available APR pipelines – we expect similar vulnerabilities to manifest.
Construct Validity. We define attack success in terms of patch generation, detection bypass, and the presence of malicious behavior. These metrics are practical, but may not fully capture downstream impact, e.g., exploitation feasibility or user-facing harm.
@@ 9. Related Work
9.1. Automated Program Repair (APR)
@@
Benchmark suites like SWE-bench and SWE-bench Multimodal assess the functional correctness of LLM-generated patches (Jimenez et al., 2024; Yang et al., 2024b), but overlook whether those patches preserve security properties or introduce vulnerabilities. While security risks of LLMs are well documented (e.g., prompt injection, unsafe code (Lyu et al., 2024)), they remain largely unexplored in LLM-based APR.
Our work fills this gap by investigating the security risks of LLM APR systems, demonstrating that adversarial bug reports can reliably induce unsafe behavior, even in state-of-the-art LLM agents.
@@ 9.2.1. Attacks Against LLM Applications
The OWASP Top 10 for LLM Applications (Wilson and Dawson, 2025) highlights common attacks and issues occurring in deployed LLM applications. We ground our research on this established industry standard to make our results applicable and relevant for real-world APR systems.
@@
We decided to not employ Training Data Poisoning (LLM03) or utilize Supply Chain Vulnerabilities (LLM05), do not exploit Insecure Plugin Design (LLM07), or Model Theft (LLM10).
@@ 9.2.2. Defenses for LLM-based Applications
We group defenses info measures that try to prevent Prompt Injection attacks, and measures that try to minimize the impact of a successful attack.
@@ 10. Conclusion
Our empirical study shows that these adversarial inputs can systematically subvert even state-of-the-art APR pipelines: in our demonstration, over 90% of the crafted bug reports resulted in patches aligned with the attacker’s objective, including CVE reversions, vulnerability injection, and CI/CD abuse. Pre- and post-APR defenses, while helpful, often failed silently (best filter only blocked 47% issues) or demanded high levels of manual effort and scrutiny (Co Pilot supported review 58%). Project test suites only slightly improve attack detection. The fundamental asymmetry—low-cost attack versus high-cost defense—further compounds the challenge.